OpenStudy (anonymous):

consider the quadratic function f(x)=2x^2 - 8x - 5 ..the smallest zero of the function is??? and how to find?? thanks!

OpenStudy (anonymous):

+ last year, a music theatre charged 60$ admission and at that price an average of 200 seats were sold for each show. a survey predicts that for every $5 increase in ticket price, ten fewer people would be expected to attend a show.. write the quadratic function to model this equation. determine the admission price that would maxinmize revenue and how many seats would be empty when revenue is maximized compared to last year's average? please help!

OpenStudy (campbell_st):

you can either complete the square of use the general quadratic formula my choice is GQF...

OpenStudy (anonymous):

how..? is this for the first question?

OpenStudy (anonymous):

the answer is -.55 but i dont know how to get it

OpenStudy (anonymous):

\[2x^{2}-8x-5\] a=2 b=-8 c=-5 Input those into the quadratic equation.

OpenStudy (campbell_st):

using my CAS calculator I got \[x= \frac{\sqrt{26} -4}{2}\]

OpenStudy (campbell_st):

thats the smaller value the larger is \[x = \frac{\sqrt{26} +4}{2}\]

OpenStudy (anonymous):

first one should be 4 - the root, not the other way around.